OpenStudy (anonymous):

Find the most general antiderivative of the function. f(x)=3:sqrt(x^2)+xsqrt(x)

OpenStudy (anonymous):

what is the colon for

OpenStudy (anonymous):

its the third root.. I wasn't sure how to write that lol

hartnn (hartnn):

\(\large f(x)=\sqrt[3]{x^2}+x \sqrt x\)

OpenStudy (anonymous):

^ exactly.

OpenStudy (anonymous):

\[\int\limits\limits_{}^{}\sqrt[3]{x^2}+x \sqrt{x} \space dx\]

hartnn (hartnn):

\(\huge \int x^ndx= \frac{x^{n+1}}{n+1}+c\)

OpenStudy (anonymous):

yes, thank you.

hartnn (hartnn):

\(\huge \int\limits_{}^{}x^{2 \over 3} + x^{3 \over 2} \space dx\)

hartnn (hartnn):

right @ChmE ?

hartnn (hartnn):

\(\huge \sqrt[3]{x^2}= (x^2)^{1/3}=x^{2/3}\)

hartnn (hartnn):

\(\huge x^1\sqrt x=x^{1+1/2}=x^{3/2}\)

OpenStudy (anonymous):

ya you're right @hartnn I did make a mistake. Sorry @monroe17

OpenStudy (anonymous):

Just another example of why you should always double check people's responses for yourself and not just take their word for it

OpenStudy (anonymous):

Yeah, I got it though. Thanks